### v1

# DeFine: Enhancing LLM Decision-Making with Factor Profiles and Analogical Reasoning

Abstract
LLMs are ideal for decision-making due to their ability to reason over long contexts and identify critical factors. However, challenges arise when processing transcripts of spoken speech describing complex scenarios. These transcripts often contain ungrammatical or incomplete sentences, repetitions, hedging, and vagueness. For example, during a company’s earnings call, an executive might project a positive revenue outlook to reassure investors, despite significant uncertainty regarding future earnings. It is crucial for LLMs to incorporate this uncertainty systematically when making decisions. In this paper, we introduce DeFine, a new framework that constructs probabilistic factor profiles from complex scenarios. DeFine then integrates these profiles with analogical reasoning, leveraging insights from similar past experiences to guide LLMs in making critical decisions in novel situations. Our framework separates the tasks of quantifying uncertainty in complex scenarios and incorporating it into LLM decision-making. This approach is particularly useful in fields such as medical consultations, negotiations, and political debates, where making decisions under uncertainty is vital.
1 Introduction
Large language models are increasingly utilized for decision-making, thanks to their advanced reasoning abilities (Eigner & Händler, 2024). While research has examined various types of reasoning, e.g., deductive, inductive, mathematical, and multi-hop reasoning, most studies have tackled simpler tasks, such as natural language inference and math word problems (Bostrom et al., 2022; Huang & Chang, 2023; Sprague et al., 2024; Mondorf & Plank, 2024). There is a significant gap in handling complex, real-world scenarios, such as making financial investment decisions (Keith & Stent, 2019), where the stakes are high and poor decisions can result in severe consequences. Therefore, it is crucial to understand how LLMs make decisions, allowing domain experts to collaborate with them to make informed, rational decisions in complex situations.
The challenges are compounded when LLMs are required to handle long contexts, extract multiple relevant pieces of information, and make decisions based on this data (Krishna et al., 2023; Laban et al., 2024). Key issues include a tendency to prioritize information at the beginning and end of the context (recency bias; Liu et al. 2023), handling inconsistencies, and mitigating hallucinations in numerical data (Hu et al., 2024a; b). Current tools such as chain-of-thought (CoT; Wei et al. 2023), tree-of-thought (ToT; Yao et al. 2023), Reflexion (Shinn et al., 2023), are designed to provide reasoning traces for LLM decisions; however, their explanations remain ambiguous. They lack precise, quantitative insights into key factors and the degree of uncertainty involved. Consequently, decision-makers are left with doubts about the reliability of these decisions and how they can be improved. There is a pressing need to enhance the verifiability of LLMs in complex decision-making scenarios to ensure their dependability and effectiveness.
We present DeFine, a new framework designed to build probabilistic factor profiles from transcripts of spoken speech that describe complex scenarios. These transcripts are often excessively long, containing ungrammatical sentences, repetitions, hedging, and vagueness (Sawhney et al., 2020; Medya et al., 2022). For example, during a quarterly earnings call, a company executive might project a positive revenue outlook to boost investor confidence, despite significant uncertainties surrounding these projections (Mukherjee et al., 2022). DeFine constructs a factor profile for each transcript that summarizes essential information into a set of factors and estimates the probabilities of potential outcomes for these factors. Moreover, we employ the Bradley-Terry model (Bradley & Terry, 1952) to identify dominant factors and evaluate how these factors collectively impact decision-making. Our factor profiles are designed to capture the nuances in spoken transcripts. This includes not just what is explicitly stated, but also the implications of details that are omitted, providing a viable method for quantifying uncertainty.
Our research integrates probabilistic factor profiles with analogical reasoning, a type of reasoning that identifies connections between similar situations to facilitate knowledge transfer from a familiar context to a novel one (Webb et al., 2023; Yasunaga et al., 2024). It helps decision-makers draw parallels between current situations and past experiences, thus effectively leveraging historical insights to inform their decisions. Instead of relying on traditional text matching, we use factor profiles to retrieve analogous examples, which identifies historical cases with similar levels of uncertainty across key dimensions. Analogical reasoning further sets our work apart from traditional Bayesian inference frameworks used in decision-making (Halawi et al., 2024; Lin et al., 2024; Liu et al., 2024). The latter often require extensive sampling during inference, which tends to increase inference costs and potentially leads to latency issues. The contributions of our research are summarized as follows.
  • We introduce DeFine, a new framework designed to enhance decision-making in LLMs. DeFine utilizes probabilistic factor profiles to quantify uncertainty in complex scenarios, along with analogical reasoning that leverages insights from similar past experiences to guide LLMs in making crucial decisions in novel situations. Our framework aims to boost the utility of LLMs by adding a layer of transparency to their decision-making processes.
  • Our research has produced actionable insights for predicting stock movement trends by analyzing earnings call transcripts. These insights enable investors to make informed, data-driven decisions. Furthermore, our approach has potential applications beyond finance, including in fields such as medical consultations and political debates (Lehman et al., 2022), where the discussions involve complex issues and the decisions made can have significant consequences. [^1] We plan to make the data and code publicly available upon acceptance to facilitate research in this area.
Figure 1: An excerpt from a typical earnings call transcript and its associated factor profile.
2 Overview and Factor Analysis
LLMs have shown great promise in financial services (Reddy et al., 2024; Lee et al., 2024), yet the specifics of their decision-making processes remain largely unexplored. In this paper, we investigate how LLMs can guide investment decisions by analyzing earnings call transcripts. The transcripts contain a rich mix of textual and numerical data, presenting a unique challenge for LLMs in making rational decisions. An earnings call is a teleconference in which the management of a public company discusses its financial results with analysts and investors for a specific reporting period, such as a quarter or a fiscal year. [^2] https://en.wikipedia.org/wiki/Earnings_call The transcript generally consists of two sections: the initial prepared remarks from the company’s executives, and a subsequent Q&A session. During these calls, executives provide a deep dive into the company’s financials, discuss key performance indicators, and share strategic plans for the future. An excerpt from a typical earnings call transcript is shown in Figure 1.
Our goal is to provide investment recommendations based on earnings call transcripts (ECTs), using a five-tier system: Strong Buy, Buy, Hold, Sell, and Strong Sell. We choose this approach over a simple binary classification (Ni et al., 2024) to give a clearer and more nuanced assessment of the investment opportunity. We highlight the key drivers in decision-making by identifying a small set of factors from a lengthy transcript (Eigner & Händler, 2024; Feng et al., 2024). For example, in an earnings call, discussions of financials such as revenue, expenses, and profit margins can be overwhelming. A factor profile helps to distill these discussions into multiple variables, effectively reducing information redundancy and allowing decision-makers to focus on the most impactful factors. Crucially, a factor profile offers a comprehensive view on an earnings call. If critical elements such as debt levels are not addressed by company executives, they can be marked as ‘unknown or uncertain.’ This contrasts with textual summaries of the transcript (Cho et al., 2021; Khatuya et al., 2024), which may be biased toward the topics emphasized by executives and discussed during the Q&A session.
2.1 Factor Profile
[…]
1. Economic Health 2. Market Sentiment and Investor Psychology 3. Political Events and Government Policies 4. Natural Disasters and Black Swan Events 5. Geopolitical Issues 6. Mergers and Major Acquisitions 7. Regulatory Changes and Legal Issues 8. Financial Health 9. Company Growth 10. Company Product Launches 11. Supply Chain 12. Technological Innovation 13. Historical Earnings Per Share (EPS) 14. Historical Revenue 15. Historical Stock Prices
Table 1: A curated set of 15 factors for forecasting stock movements following earnings.
In this study, we focus on a curated set of 15 factors, categorized into three groups: macroeconomic influences (e.g., economic health, market sentiment), company-specific dynamics (e.g., mergers and major acquisitions, product launches), and historical financial metrics (e.g., past earnings, stock prices). These factors were carefully selected through an iterative process of querying the LLM for key variables crucial in forecasting stock movements following earnings announcements. We intentionally limited our variable set to 15 factors, each with two to three potential outcomes, as detailed below. By distilling the analysis to a few significant predictors, our approach balances complexity and performance, while also allowing for future integration of domain-specific factors identified by financial analysts.
  • Macroeconomic Influences. These encompass broad economic factors that affect the entire market or large segments of it. This includes the overall economic health, market sentiment, political events, natural disasters and geopolitical issues (Liu et al., 2024). Each factor leads to two potential outcomes; for instance, natural disasters might cause a ‘Major Impact’ by disrupting economies and global supply chains, and directly affecting market performance; the ‘Unknown or Uncertain’ outcome reflects the unpredictability of such events.
  • Company-Specific Dynamics. These factors are linked to the internal operations and strategic decisions of individual companies, such as mergers and acquisitions, regulatory changes, financial health, company growth potential, product launches, and issues within the supply chain. Each factor can result in one of two potential outcomes. For example, a ‘Positive Outlook’ on regulatory changes can open up new business opportunities, whereas ‘Unknown or Uncertain’ could signify regulatory uncertainties that lead to financial challenges.
  • Historical Financial Metrics. Important metrics include historical earnings per share (EPS), revenue trends, and past stock price movements. Each factor can result in three outcomes: ‘Bullish’, where metrics like earnings per share, revenue, and stock prices consistently rise, indicating strong financial health; ‘Stable’, characterized by steady movements; ‘Bearish,’ marked by declining financial figures, possibly leading investors to be pessimistic about the company’s future performance.
We make use of the structured output capability of GPT-4o-2024-08-06 to extract factor profiles from earnings call transcripts. Following the framework set by Liu et al. (2024), we provide the LLM with a list of factors, their potential outcomes, and associated verbalized likelihoods. For each factor, the analysis involves two steps: first, the LLM creates a concise summary specific to that factor from the transcript; second, it assigns a verbalized likelihood to each possible outcome, ranging from “very unlikely” to “very likely.” Specifically, the likelihoods of outcomes, such as EPS, revenue trends, and historical stock prices, are derived from the company’s historical financial data. An example of the factor profile is shown in Figure 1, and the prompts used are detailed in the Appendix.
To convert these categorical likelihoods into probabilities, we employ the following normalization process: let $P_{i,j}$ denote the likelihood associated with the $j$-th outcome for the $i$-th factor. Here, verbalized likelihoods are converted to numerical values using the mapping {very unlikely=1, unlikely=2, somewhat unlikely=3, somewhat likely=4, likely=5, very likely=6}. Then, the probability $P(O_{ij}|X)$ is calculated as $P(O_{ij}|X)=\frac{P_{i,j}}{\sum_{k}P_{i,k}}$, ensuring the sum of outcomes for each factor equals 1. Alternative techniques, such as instructing the LLM to “distribute 10 points among the outcomes”, have been explored (Yang et al., 2024), our initial evaluation reveals that using verbalized likelihoods followed by normalization improves prediction accuracy compared to these direct probability distribution methods.
2.2 Analyzing Key Factors Using the Bradley-Terry Model
The Bradley-Terry model is a probabilistic framework used for estimating the relative strengths of items based on pairwise comparisons, and the outcome of each comparison indicates which of the two items is ‘better’ in a specific context (Bradley & Terry, 1952). This model has been widely used for ranking purposes in sports tournaments, LLM preference studies, and other domains where pairwise comparison data is available (Hu et al., 2023; Zhu et al., 2024). In this model, we estimate parameters that represent the strength of each factor. These parameters are generally presented on a logistic scale, where the probability that factor A is considered more significant than factor B is modeled as:
$P(A>B)=\frac{e^{\beta_{A}}}{e^{\beta_{A}}+e^{\beta_{B}}}$ (1)
Here, $\beta_{A}$ and $\beta_{B}$ represent the strengths of factors A and B, respectively. The estimated parameters are often exponentiated, so that $p_{i}=e^{\beta_{i}}$ measures the relative strength of each factor. A higher value indicates a stronger influence. In determining which factors to prioritize in a post-earnings analysis, those with higher Bradley-Terry scores are considered more crucial.
Consider a comparative analysis of two earnings call transcripts, A and B, transcript A is more likely to lead to favorable stock movements than transcript B (A $\succ$ B). We obtain such pairwise comparisons based on target labels; with ‘strong-buy’ ranked higher than ‘hold’, ‘sell’, and ‘strong-sell’; ‘buy’ outranking ‘sell’ and ‘strong-sell’; and ‘hold’ surpassing ‘strong-sell’. The comparison of A and B will involve creating a set of factor-outcome pairwise comparisons, where each outcome in transcript A is preferable to that in transcript B: $O_{\cdot,\cdot}^{(A)}\succ O_{\cdot,\cdot}^{(B)}$, suggesting that the factors associated with transcript A outperform those in transcript B.
We further consider the weight-adjusted effect of comparisons between factors. Our method compares the influence of factors from transcripts A and B by calculating an ‘expected occurrence’, which is determined by multiplying the likelihood of these factors appearing in both transcripts, $P(O_{ij}|X^{(A)})\times P(O_{ij}|X^{(B)})$. This approach provides a probability-based comparison, offering a more detailed evaluation than simple counting methods. These expected occurrences then feed into a Bradley-Terry model matrix $W$. The model helps to estimate the relative importance of each factor by assigning a coefficient $p_{x}$ to each outcome $O_{ij}$, indicating its influence on stock investment decisions. We refine these estimates using an EM-like algorithm, which iteratively adjusts and normalizes $p_{x}$ to best fit the observed data.
$p^{\prime}_{x}=W_{x}\left(\sum_{y\neq x}\frac{w_{xy}+w_{yx}}{p_{x}+p_{y}}\right)^{-1}\quad p_{x}=\frac{p^{\prime}_{x}}{\sum_{y=1}^{M}p^{\prime}_{y}}$ (2)
3 Bayesian Decision-Making
In Bayesian decision-making, utility functions play a crucial role in navigating uncertainty (Halawi et al., 2024; Lin et al., 2024; Ye et al., 2024). A Bayesian framework updates beliefs about possible outcomes. Decisions are then made by evaluating the expected utility for each possible action, which involves calculating the utility across the updated beliefs. This method ensures that choices are made to maximize expected utility, so decisions are aligned with the decision-maker’s preferences and risk tolerance.
Concretely, to compute $P(O_{ij}|X)$, we construct a probabilistic factor profile from a given earnings call transcript, where $O_{ij}$ represents the $j$-th outcome of the $i$-th factor. The likelihood $P(Y|O_{ij})$, which estimates how the $j$-th outcome influences stock investment decisions, is calculated using the Bradley-Terry model. This model provides a framework for quantifying the impact each factor outcome has on the decision-making process. Using these probabilities, the Bayesian decision-making formula integrates over all factors and their potential outcomes to determine the optimal action. The overall decision is derived by:
$\hat{Y}=\arg\max_{Y}\sum_{i}\sum_{j}P(Y|O_{ij})P(O_{ij}|X)$ (3)
The parameters calculated by the Bradley-Terry model for $P(Y|O_{ij})$ help us determine how each factor influences stock movements. During our testing phase, transcripts are assigned to one of five decision categories based on their computed scores. For example, if the ground truth indicates there are $k$ ‘strong buy’ recommendations, the top $k$ scoring transcripts are classified correspondingly as ‘strong buy’. This approach uses probabilistic factor profiles in conjunction with Bradley-Terry modeling to identify influential factors, providing a transparent method for understanding decision-driving elements. Moving forward, we extend beyond individual factors by examining analogous cases that directly influence decisions.
4 Analogical Reasoning
Analogical reasoning, which involves drawing parallels between similar situations (Webb et al., 2023; Ozturkler et al., 2023; Yuan et al., 2024; Sourati et al., 2024; Yasunaga et al., 2024), is an effective method for decision-making. This approach is particularly useful when analyzing how stocks react to earnings announcements by referencing past, similar events. For example, in the tech sector, stocks often show high volatility after earnings calls that introduce significant technological updates, even if the revenue and EPS meet expectations. If a tech company is rumored to discuss a new technology trend in its upcoming earnings announcement, using this method, we can infer that this company’s stock might also experience increased volatility. Investors might use this analysis to make investment decisions or hedge against potential volatility.
Accurately identifying analogous examples from earnings call transcripts is crucial. We propose a method that utilizes probabilistic factor profiles, denoted as $P(O_{ij}|X)$, where $O_{ij}$ represents the $j$-th outcome of the $i$-th factor. To measure the similarity between profiles, we calculate the Kullback-Leibler (KL) divergence, which quantifies the information loss when one probability distribution approximates another. The KL divergence is computed as follows:
$D_{KL}(P||Q)=\sum_{i=1}^{n}\sum_{j=1}^{m}P(O_{ij}|X)\log\frac{P(O_{ij}|X)}{Q(O_{ij}|X_{c})}$ (4)
Here, $P$ represents the factor profile for the target transcript, and $Q$ denotes the profile for a comparative transcript $X_{c}$ from our training set. Transcripts with lower KL divergence values are considered more analogous, and therefore more likely to influence investor decisions similarly.
During testing, we identify the Top-K profiles that show the least divergence from a test instance’s profile and present these as analogical examples for the LLM to consider when reasoning about stock movements. The LLM is asked to select the most analogous example from the Top-K and carefully evaluates the current test instance to make its prediction. This approach ensures that the alignment between profiles is contextually appropriate, thereby drawing meaningful comparisons across different transcripts. By focusing on factor profiles rather than full transcripts or their summaries, we emphasize key market-moving information, avoiding unnecessary details. For example, Google and Broadcom could have analogous profiles even though their discussions in earnings calls might vary widely. Using factor profiles as analogous examples also requires significantly fewer tokens within the context window than full transcripts would.
Data Statistics Num. of Transcripts 11,950 Num. of Companies 869 Avg. #Tokens per Transcript 10,187 Avg. #QA Pairs per Transcript 10 Avg. #Trans per Company 14 Avg. #Speakers per Transcript 12 Year Range 2017–2024
Table 2: Our dataset includes 11,950 earnings call transcripts from 800+ companies.
5 Data Collection
Our dataset contains 11,950 earnings call transcripts from S&P 500 and NASDAQ 500 companies, gathered from the Motley Fool over the period of 2017–2024. The Motley Fool is a well-regarded financial service website that regularly publishes earnings call transcripts from U.S. companies. We make sure to follow their terms of use carefully during data collection. We do not use audio recordings or analyze acoustic or prosodic features. Each transcript is formatted as a JSON object, including the company’s stock ticker, the date of the earnings announcement, participant names and their affiliations, executive prepared remarks, and a series of question-answer pairs from the Q&A session. Table 2 presents the statistics of our dataset. Each transcript averages 10,187 tokens and 133 sentences. They are sourced from 869 companies, each contributing an average of 14 transcripts. We obtain company stock prices from Yahoo Finance via the yfinance package and financial metrics such as revenue and earnings per share (EPS) from Alpha Advantage. Our dataset spans from 2017 to 2024. It enhances previous studies which examined earnings call transcripts from 2002–2010 (Li et al., 2020); these earlier transcripts may already be used in LLM pretraining. To avoid data contamination, we established a new test set consisting of the most recent 587 transcripts from 2024, which are beyond the pretraining cut-off date for LLMs.
We seek to make stock investment decisions by analyzing earnings call transcripts and focusing on performance over the 30-day period. We establish the ground truth decision on the 30th day following each earnings announcement (Sonkiya et al., 2021): a stock drop exceeding 5% corresponds to a ‘strong sell’ decision, a decrease between 2% and 5% leads to a ‘sell’, fluctuation within -2% to +2% is labeled ‘hold’, an increase between 2% and 5% is labeled a ‘buy’, and an increase above 5% is a ‘strong buy’. In our test set, the distribution of these labels is as follows: ‘strong buy’ at 34%, ‘buy’ at 15%, ‘hold’ at 21%, ‘sell’ at 9%, and ‘strong sell’ at 21%. This distribution is generally balanced, reflecting a slightly bullish market trend in 2024.
6 Experiments
In this section, we evaluate the decision-making performance of various systems, analyze the key factors that influence stock movement predictions, and conduct an analysis of analogical reasoning.
System Recall Prec. F1 Accu. Label Recall Prec. F1
LLM+CoT+Trans 21.56 33.66 13.52 19.59 Strong Sell 7.32 37.50 12.24
LLM+CoT+Summ 22.77 16.17 14.12 20.61 Sell 5.56 9.09 6.90
LLM+CoT+Factors 24.38 28.58 17.26 22.32 Hold 29.84 28.24 29.02
DeLLMa (Liu et al., 2024) 38.30 23.14 16.68 22.35 Buy 44.83 18.93 26.62
DeFine (Ours) 26.15 27.67 23.73 29.64 Strong Buy 43.22 44.56 43.88
Table 3: (Left) We show the accuracy and macro-averaged F-scores for various systems. Our system, DeFine, which combines factor profiles with analogical reasoning, achieves the best performance. (Right) DeFine’s performance across five categories: Strong Sell, Sell, Hold, Buy, and Strong Buy.
6.1 Decision-Making with DeFine
We test our system, DeFine, against different decision-making strategies: (a) LLM+CoT+Trans: We feed the entire earnings call transcript to the LLM and then use the chain-of-thought to assign a label, with both labels and their interpretations also provided to the LLM. (b) LLM+CoT+Summ and LLM+CoT+Factors: These approaches use a summarize-then-predict strategy. LLM+CoT+Summ simplifies the transcript into a textual summary, while LLM+CoT+Factors condenses it into a factor profile. It is considered a structured summary, unlike the textual summary produced by direct LLM prompting. Details on the prompts used for these methods can be found in the Appendix.
Our system, DeFine, utilizes analogical reasoning by analyzing five analogous cases identified using KL-divergence as the distance metric. It examines these cases alongside the current factor profile to predict an appropriate label. In contrast, DeLLMa uses a decision theory approach and has shown strong performance in agriculture planning and finance (Liu et al., 2024). For this approach, we pair each factor profile with possible labels and choose the top-ranked outcome as the final decision.
In Table 3 (left), we present the accuracy and macro-averaged F-scores for various systems, all using GPT-4o-2024-08-06. Our new system, DeFine, which combines factor profiles with analogical reasoning, achieves the best performance. It surpasses the strong baseline system, DeLLMa, which involves ranking state-action pairs based on their preference levels as determined by the LLM. We find that LLMs generally make more accurate decisions when working with summaries rather than full transcripts; those transcripts typically contain around 10k tokens. This finding underscores the complexity of extracting and weighing key factors from lengthy transcripts, a task that remains challenging for most LLMs. In contrast, our factor profile method proves advantageous as it provides a balanced view of both macroeconomic factors and company-specific details, which are essential for rational decision-making.
Figure 2: A comparison of confusion matrices from the LLM+CoT+Trans, DeLLMa, and DeFine methods. While LLM+CoT+Trans and DeLLMa lean towards ‘Buy (B),’ DeFine offers more balanced outcomes across all decision categories, showing notable improvement in ‘Strong Buy (SB),’ ‘Buy (B),’ ‘Hold (H),’ and ‘Sell (S)’ decisions.
We further analyze DeFine’s performance across five categories: Strong Sell, Sell, Hold, Buy, and Strong Buy. Results are shown in Table 3 (right). DeFine performs best at ‘Strong Buy’ recommendations and faces challenges with ‘Strong Sell’ categories. This may be due to its reliance on earnings call transcripts, which often contain optimistic remarks from executives aimed at reassuring investors, potentially skewing predictions away from ‘Strong Sell.’ Figure 2 includes a comparison of confusion matrices from the LLM+CoT+Trans, DeLLMa, and DeFine methods. While LLM+CoT+Trans and DeLLMa predominantly lean towards ‘Buy,’ DeFine offers more balanced outcomes across all decision categories, showing notable improvement in ‘Strong Buy,’ ‘Buy,’ ‘Hold,’ and ‘Sell’ decisions.
6.2 Influential Factors
We develop three variations of our DeFine-BT approach, each using the Bradley-Terry model for pairwise comparisons in different contexts: DeFine-BT-Same Sector compares companies within the same sector, DeFine-BT-Cross Sectors examines companies across different sectors, and DeFine-BT-Same Company analyzes a company’s current earnings call transcript against its historical ones. To ensure fairness, we maintain the same number of pairwise comparisons across all three settings, downsampling where necessary. According to the F-scores presented in Table 4, all DeFine-BT variants outperform both the random baseline, which assigns investment decisions randomly from five possible labels, and DeLLMa on the test set.
Among the three variants, DeFine-BT-Cross Sector achieves the highest scores in both F-Score and Accuracy. This indicates that considering pairwise comparisons between earnings announcements from a diverse range of companies can enhance predictions of stock movements. Table 7 illustrates the performance of DeFine-BT-Cross Sector, which was trained on one sector and tested on another. For this analysis, 100 earnings call transcripts were selected from each of the 11 financial sectors: Technology, Healthcare, Financial Services, Consumer Defensive, Energy, Industrials, Utilities, Basic Materials, Real Estate, Consumer Cyclical, and Communication Services.
F1 Accu. Random Baseline 18.00 19.11 DeLLMa ((Liu et al., 2024)) 16.68 22.53 DeFine-BT-Same Sector 20.11 22.15 DeFine-BT-Same Company 20.42 23.68 DeFine-BT-Cross Sectors 24.45 27.43
Table 4: Among the three variants, DeFine-BT-Cross Sector achieves the highest scores, suggesting that considering pairwise comparisons from a diverse range of companies can enhance the predictions of stock movements.
Factor (Outcome) Salience - Regulatory changes and legal issues 0.0364 happened (positive outlook) - Natural disasters and other black 0.0360 swan events (major impact) - Political events and government 0.0349 policies (major upheaval) - Geopolitical issues (escalation to 0.0345 conflict) - Supply chain (positive outlook) 0.0322 - Tech innovation (positive outlook) 0.0317 - Historical stock price change (bullish) 0.0316 - Historical EPS (bullish) 0.0315 - Financial health (positive outlook) 0.0311
Table 5: Influential factors and outcomes that drive bullish investment decisions in the Consumer Defensive sector, such as food and beverage, household products, and grocery stores.
Factor (Outcome) Salience - Economic health (unknown or uncertain) 0.0362 - Market sentiment and investor 0.0350 psychology (unknown or uncertain) - Company growth (unknown or uncertain) 0.0338 - Supply chain (unknown or uncertain) 0.0326 - Geopolitical issues (escalation to conflict) 0.0322 - Historical revenue (decline) 0.0319 - Historical stock price change (bullish) 0.0318 - Tech innovation (unknown or uncertain) 0.0315 - Natural disasters and other black 0.0315 swan events (major impact) - Political events and government 0.0313 policies (major upheaval)
Table 6: Factors and outcomes that drive bullish investment decisions in the Technology sector, including industry leaders such as Apple, Microsoft, Amazon, Google, and Meta.
Tables 6 and 6 highlight influential factors impacting investment decisions in the Consumer Defensive and Technology sectors, as identified by the Bradley-Terry model. In Consumer Defensive, which includes industries like food and beverage, household products, and grocery stores, significant drivers are natural disasters and black swan events, political events and government policies, and geopolitical issues. These challenging macroeconomic circumstances often lead to buy-in decisions from investors. In contrast, the Technology sector, with industry leaders such as Apple, Microsoft, Amazon, Google, Meta, and Nvidia, shows that decisions to invest often hinge on unclear or uncertain factors. Technology stocks have seen considerable growth from 2017–2024. This pattern suggests that investment models may favor purchases in these companies despite encountering negative issues in earnings announcements.
In Figure 4, we analyze the probability of positive and negative factor outcomes, represented as a continuous random variable, and plot its probability density function (PDF) for various investment decisions. Highlighted sections illustrate where the gaps between strong buy (red) and strong sell (blue) decisions are most pronounced. Our analysis indicates that buy decisions often occur when the probability of positive outcomes is relatively low (about 0.2-0.3) and the likelihood of negative outcomes is moderate to high (ranging from 0.3 to 0.65), but not overly negative. Conversely, sell decisions tend to occur when negative outcome probabilities are minimal (about 0.1-0.2). These observations suggest that rational investment decisions can sometimes appear counterintuitive: essentially, selling high and buying low. We find that a thorough analysis of various factors is advantageous. Our approach incorporates not just the known issues but also the uncertain or hidden factors, thereby enhancing the decision-making process.
6.3 Insights into Analogical Reasoning
Analogical reasoning utilizes a select number of analogous examples, denoted as $K$, to inform decision-making in LLMs. In Figure 3, we adjust $K$ from 3 to 9 and observe its impact on the F-Score. In these experiments, we use the majority vote from the $K$ examples as the final prediction. We find that $K=4$ achieves the highest performance, potentially due to some tie-induced randomness compared to odd numbers. Typically, odd numbers for $K$ are preferred for majority voting to avoid ties, with $K=3,5,7$ showing similar effectiveness. For our system, DeFine, we have opted for $K=5$ to strike a balance between providing enough analogous examples and maintaining a manageable context length for the LLM.
Figure 3: Analogical reasoning utilizes a number of analogous examples to inform decision-making in LLMs. We adjust $K$ from 3 to 9 and observe its impact on the F-Score.
Moreover, we examine how the most analogous examples influence DeFine’s predictions. Our study finds that in 69% of cases, the LLM’s predictions match the labels from the most analogous examples. In the other 31% of cases, the LLM chooses to make its own predictions. E.g., when the analogous example is labeled “Strong Buy,” DeFine concurs with “Strong Buy” in 63% of cases. It opts for “Buy” in 26% and “Hold” in 11% of the cases. Conversely, when the example is “Strong Sell,” DeFine agrees with “Strong Sell” 50% of the time, chooses “Sell” in 25% of cases, and “Hold” in 12.5%. These results indicate that while DeFine effectively utilizes analogous historical data to inform its predictions, it also critically evaluates the current factor profiles, demonstrating a balanced approach in its decision-making abilities.
Figure 4: We analyze and plot the probability density function (PDF) of positive and negative factor outcomes for different investment decisions. Highlighted sections illustrate where the gaps between strong buy (red) and strong sell (blue) decisions are most pronounced.
7 Related Work
Analogical Reasoning. This type of reasoning identifies connections between similar, though not identical, situations to transfer knowledge from a known context to a new one (Webb et al., 2023; Ozturkler et al., 2023; Yu et al., 2024; Yuan et al., 2024; Sourati et al., 2024; Yasunaga et al., 2024). It helps decision makers draw parallels between current situations and past experiences, effectively leveraging historical insights. Analogical reasoning plays a crucial role in various fields, e.g., doctors apply knowledge from one disease to diagnose another, and lawyers use past rulings to argue new cases (Lehman et al., 2022; Charmet et al., 2022; Cao et al., 2024a). This ability to recognize and use similarities in different situations is important for decision-making.
While zero-shot analogical reasoning is a desired capability for LLMs, recent studies show they lack the robustness and generality of human analogy-making, as evidenced by counterexamples in tasks such as letter string analogies (Hodel & West, 2024; Lewis & Mitchell, 2024). Musker et al. (2024) test both humans and LLMs on tasks that require transferring semantic structure and content between domains. Yasunaga et al. (2024) introduce analogical prompting, where LLMs self-generate relevant examples using prompts such as “# Recall relevant problems and solutions:” before solving the original problem; Qin et al. (2024) find that the accuracy of self-generated examples is key to eliciting such capability. Unlike previous research, our study employs probabilistic factor profiles to model analogical reasoning, grounding our approach in solid mathematical principles.
Tech FS Health CC Ind CS CD Energy RE BM Util Technology (Tech) 15.40 17.99 17.39 12.10 13.15 18.19 25.85 27.09 13.82 22.86 26.67 Financial Services (FS) 15.96 17.96 26.84 7.99 10.21 26.22 15.45 4.80 13.37 21.37 0.00 Healthcare (Health) 16.73 19.80 17.89 21.85 28.46 10.86 20.23 18.73 3.64 43.45 73.33 Consumer Cyclical (CC) 18.14 11.02 19.38 19.39 15.86 9.49 17.70 17.40 12.19 22.22 36.67 Industrials (Ind) 17.02 11.14 14.37 11.24 18.81 15.93 19.48 25.11 3.20 24.44 0.00 Communication Services (CS) 18.61 14.68 18.87 14.03 19.47 33.70 10.71 16.99 11.87 10.26 13.33 Consumer Defensive (CD) 24.91 21.71 19.15 19.89 21.38 2.67 23.09 12.50 9.72 29.52 50.00 Energy 19.49 16.50 23.62 14.25 19.03 8.90 19.03 8.98 12.10 28.79 0.00 Real Estate (RE) 22.86 15.74 16.76 14.08 12.34 4.00 15.61 11.28 12.34 43.18 0.00 Basic Materials (BM) 20.67 13.69 15.26 18.18 29.64 9.52 21.19 17.19 17.10 16.67 37.50 Utilities (Util) 17.82 27.75 23.15 26.25 12.61 25.49 20.63 5.70 18.40 14.29 53.33
Table 7: The performance of DeFine-BT was evaluated by training it on one financial sector and testing it on another using 100 earnings call transcripts from each of the 11 sectors.
LLM Decision-Making under Uncertainty. The use of LLMs in decision-making has surged due to their remarkable ability to reason over complex scenarios (Halawi et al., 2024; Lin et al., 2024; Ye et al., 2024; Band et al., 2024). However, the challenge of balancing a multitude of often conflicting factors in decision making remains understudied. For example, Falck et al. (2024) investigate whether adding more data points in in-context learning reduces uncertainty, as typically expected in Bayesian learning, and find evidence against this theory. The DeLLMa framework (Liu et al., 2024) incorporates uncertainty into LLM decision-making using Bayesian networks and has been tested on tasks such as agriculture planning and finance. Feng et al. (2024) employ LLM entailment to map factors to context and utilize trained Bayesian models for probability estimation. Our work builds on these initiatives by integrating analogical reasoning with factor profiles to enhance the accuracy and transparency of LLM decision-making.
Financial Forecasting. Recent advancements in LLMs have revolutionized traditional financial tasks (Keith & Stent, 2019; Sawhney et al., 2020; 2021; Chuang & Yang, 2022; Ang & Lim, 2022; Sang & Bao, 2022; Medya et al., 2022; Wang et al., 2023; Koa et al., 2024; Srivastava et al., 2024). Notably, Chen et al. (2022) introduce FinQA, a dataset constructed from financial statements for assessing LLMs’ multi-step numerical reasoning. Moreover, TAT-QA (Zhu et al., 2021) tackles QA over tabular and textual data; FiNER (Loukas et al., 2022) focuses on numerical entity recognition; DocFinQA (Reddy et al., 2024) is a dataset designed for long-document financial QA; RiskLabs (Cao et al., 2024b) employs LLMs for financial risk assessments. Nie et al. (2024) provide a comprehensive survey on the use of LLMs across various financial domains. Our study focuses on analyzing earnings transcripts to understand how LLMs handle the ambiguities inherent in spoken language, thus providing insight into their decision-making under uncertainty. The research findings have broader applications including medical consultations, negotiations, and political debates.
8 Conclusion
We propose DeFine, a new framework for decision-making in complex scenarios, such as those encountered in corporate earnings calls. By combining probabilistic factor profiles with analogical reasoning, this framework not only captures the uncertainties embedded in earnings call transcripts but also allows the LLM to apply previous insights to new challenges more efficiently. Our approach surpasses strong baseline models and enhances the practical utility of LLMs by identifying analogous examples. The DeFine framework offers a promising avenue for navigating complex data and supporting decision-making processes.
9 Limitations
The effectiveness of the DeFine framework, as presented in this paper, is predominantly based on controlled experimental conditions. While the framework has been designed to enhance decision-making capabilities through the use of probabilistic factor profiles and analogical reasoning, actual outcomes may vary when applied in real-world scenarios. Users should be aware that the framework’s performance can be influenced by various external factors including data quality, context-specific nuances, and the dynamic nature of real-world environments. We encourage users to consider these variables when implementing and adapting the DeFine approach to ensure its optimal application and to mitigate potential discrepancies between expected and actual results.
Appendix A Appendix
[…]
PromptBox s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark, colback=light!30!white, fonttitle=, fontupper=, title=Constructing a Factor Profile from an Earnings Call Transcript,
{PromptBox}
System Message
You are a financial analyst specializing in earnings call transcripts. You will receive the complete transcript of an earnings call, which includes both the prepared remarks and the Q&A session. Your job is to identify the key factors from the transcript and assign probabilities to the potential outcomes of these factors.
User Message
Your task is to conduct a comprehensive analysis of the earnings call transcript below. Be sure to accurately capture the important factors and estimate the likelihood of each factor resulting in specific outcomes.
Earnings Call Transcript for Company {Company}
# Prepared Remarks
Speaker: {Speech…}
# Questions and Answers
Analyst: {Speech…}
Please analyze the above earnings call transcript, focusing on the following key factors:
{Enumerate factors, descriptions, and outcomes}
1. Economic health: Economic health refers to the overall stability and performance of the economy, reflected in factors like growth, employment, inflation, and market confidence. Outcomes: {positive-outlook, unknown-or-uncertain}
2. Market sentiment and investor psychology: Market sentiment reflects the overall mood or attitude of investors toward a particular market, influenced by news, economic data, and global events. Investor psychology refers to the emotions and cognitive biases that drive decisions, often leading to behaviors like fear-driven selling or greed-fueled buying. Outcomes: {optimistic, unknown-or-uncertain}
Please take the time to thoroughly understand the transcript. For each key factor, provide a detailed summary based on the given transcript. Then, review all associated outcomes and assess the likelihood of each outcome. The likelihood should be strictly selected from the following options: {very likely, likely, somewhat likely, somewhat unlikely, unlikely, very unlikely}. Format your response in JSON.
# Example Output:
{JSON output example}
# Your Output:
Figure 5: Constructing a Factor Profile from an Earnings Call Transcript
[…]
histPrompt s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark,colback=light!30!white, fonttitle=, fontupper=, title=A Prompt to Analyze Trends Based on Historical Financial Metrics,
[…]
ARPrompt s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark,colback=light!30!white, fonttitle=, fontupper=, title=Using Analogical Reasoning to Make Investment Decisions,
{ARPrompt}
System Message
You’re a financial analyst who specializes in giving investors buy or sell recommendations by thoroughly analyzing earnings call transcripts.
User Message
Here are several example company profiles. Each profile highlights key factors from an earnings call transcript and probabilities for potential outcomes based on those factors. Each profile represents a specific company and is based on its historical earnings call data. Your job is to pick the most analogous example and use its strategy to solve the initial problem.
Example Company Profile 1:
{Factor Profile 1}
Analyst recommendation: {Action 1}
Example Company Profile 2:
{Factor Profile 2}
Analyst recommendation: {Action 2}
Example Company Profile 3:
{Factor Profile 3}
Analyst recommendation: {Action 3}
Example Company Profile 4:
{Factor Profile 4}
Analyst recommendation: {Action 4}
Example Company Profile 5:
{Factor Profile 5}
Analyst recommendation: {Action 5}
**Initial Problem**
Based on your analysis of the earnings call for {Company Name} held on {Announcement Date}, decide on the most likely analyst recommendation for the next 30 days from these options:
- Action 1: strong buy: The stock price will increase by more than 5%
 - Action 2: buy: The stock price will increase by 2% to 5%
 - Action 3: hold: The stock price is expected to remain stable, fluctuating between -2% to 2%
 - Action 4: sell: The stock price will decrease by 2% to 5%
 - Action 5: strong sell: The stock price will decrease by more than 5%
Below is the company profile summarized from {Company Name}’s earnings call on {Announcement Date} and the historical price trend probabilities judged by an analyst:
{Factor Profile Constructed Using an Earnings Call Transcript}
**Solve the Initial Problem**
Please respond with the analyst recommendation for this stock in JSON format, including these keys: (‘idx’, ‘recommendation’, ‘justification’). ‘idx’ is the index of the most analogous example profile, and ‘recommendation’ should be one of the actions mentioned above for 30 days of trading, and ‘justification’ should clearly explain your recommendation using the strategy you learned from the selected example company profile.
Figure 6: Using Analogical Reasoning to Make Investment Decisions
[…]
CoTPrompt s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark,colback=light!30!white, fonttitle=, fontupper=, title=A Prompt that Uses Chain-of-Thought to Make Investment Decisions,
{CoTPrompt}
System Message
You’re a financial analyst spcializing in giving investors buy or sell recommendations by thoroughly analyzing earnings call transcripts.
User Message
Based on your analysis of the earnings call for {Company Name} held on {Announcement Date}, decide on the most likely analyst recommendation for the next 30 days from these options:
 - Action 1: strong buy: The stock price will increase by more than 5%
 - Action 2: buy: The stock price will increase by 2% to 5%
 - Action 3: hold: The stock price is expected to remain stable, fluctuating between -2% to 2%
 - Action 4: sell: The stock price will decrease by 2% to 5%
 - Action 5: strong sell: The stock price will decrease by more than 5%
Below is the {Factor Profile, Transcript or Summary} from {Company Name}’s earnings call on {Announcement Date}:}
{Factor Profile, Transcripts or Summary}
Please think step by step and respond with the analyst recommendation for this stock in JSON format, including these keys: (‘thoughts’, ‘recommendation’, ‘justification’). ‘Thoughts’ should be your detailed reasoning steps, ‘recommendation’ should be one of the actions mentioned above for 30 days trading, ‘Justification’ should clearly explain your recommendation using the strategy you learned from the selected example company profile.
Figure 7: A Prompt that Uses Chain-of-Thought to Make Investment Decisions
{histPrompt}
System Message
You are a financial analyst specializing in historical data analysis, including stock prices, earnings per share (EPS), and revenue. Your goal is to assess the likelihood of different market trends based on past data.
User Message
The potential outcomes to consider are: {bullish, stable, and bearish}. For each outcome, please assign a likelihood level from the following options: {very likely, likely, somewhat likely, somewhat unlikely, unlikely, very unlikely}.
Below, you will be provided with a historical data table: {Data Name}:{Description}
{Historical Data Table}
 Date                 Close Price
 2023-07-31  195.22
 2023-08-01  195.46
{… until the date of the earnings announcement.}
Please analyze this historical data and provide the likelihood of each outcome in JSON format.
# Example Output:
{"historical EPS":{"bullish": very likely, "stable":somewhat likely, "bearish": unlikely}}
# Your Output:
Figure 8: A Prompt to Analyze Trends Based on Historical Financial Metrics

### v2

@@
-DeFine: Enhancing LLM Decision-Making with Factor Profiles and Analogical Reasoning
+DeFine: Decision-Making with Analogical Reasoning over Factor Profiles
 Abstract
-LLMs are ideal for decision-making due to their ability to reason over long contexts and identify critical factors.
-However, challenges arise when processing transcripts of spoken speech describing complex scenarios.
-These transcripts often contain ungrammatical or incomplete sentences, repetitions, hedging, and vagueness.
-For example, during a company’s earnings call, an executive might project a positive revenue outlook to reassure investors, despite significant uncertainty regarding future earnings.
+LLMs are ideal for decision-making thanks to their ability to reason over long contexts.
+However, challenges arise when processing speech transcripts that describe complex scenarios, as they are verbose and include repetition, hedging, and vagueness.
+E.g., during a company’s earnings call, an executive might project a positive revenue outlook to reassure investors, despite uncertainty regarding future earnings.
 It is crucial for LLMs to incorporate this uncertainty systematically when making decisions.
-In this paper, we introduce DeFine, a new framework that constructs probabilistic factor profiles from complex scenarios.
-DeFine then integrates these profiles with analogical reasoning, leveraging insights from similar past experiences to guide LLMs in making critical decisions in novel situations.
-Our framework separates the tasks of quantifying uncertainty in complex scenarios and incorporating it into LLM decision-making.
-This approach is particularly useful in fields such as medical consultations, negotiations, and political debates, where making decisions under uncertainty is vital.
+In this paper, we introduce DeFine, a modular framework that constructs probabilistic factor profiles from complex scenarios.
+It then integrates these profiles with analogical reasoning, leveraging insights from similar past experiences to guide LLMs in making critical decisions in new situations.
+Our framework separates the tasks of quantifying uncertainty and incorporating it into LLM decision-making.
+This approach is particularly useful in areas such as consulting and financial deliberation, where making decisions under uncertainty is vital.
+[…]
+tcb@breakable
+DeFine: Decision-Making with Analogical Reasoning over Factor Profiles
+Yebowen Hu, Xiaoyang Wang, Wenlin Yao, Yiming Lu,§ Daoan Zhang Hassan Foroosh, Dong Yu, Fei Liu§ University of Central Florida   Tencent AI Lab, Seattle University of Rochester   §Emory University {yebowen.hu, hassan.foroosh}@ucf.edu    daoan.zhang@rochester.edu {shawnxywang, wenlinyao, dyu}@global.tencent.com  fei.liu@emory.edu
 1 Introduction
-Large language models are increasingly utilized for decision-making, thanks to their advanced reasoning abilities (Eigner & Händler, 2024).
-While research has examined various types of reasoning, e.g., deductive, inductive, mathematical, and multi-hop reasoning, most studies have tackled simpler tasks, such as natural language inference and math word problems (Bostrom et al., 2022; Huang & Chang, 2023; Sprague et al., 2024; Mondorf & Plank, 2024).
-There is a significant gap in handling complex, real-world scenarios, such as making financial investment decisions (Keith & Stent, 2019), where the stakes are high and poor decisions can result in severe consequences.
-Therefore, it is crucial to understand how LLMs make decisions, allowing domain experts to collaborate with them to make informed, rational decisions in complex situations.
-The challenges are compounded when LLMs are required to handle long contexts, extract multiple relevant pieces of information, and make decisions based on this data (Krishna et al., 2023; Laban et al., 2024).
-Key issues include a tendency to prioritize information at the beginning and end of the context (recency bias; Liu et al.
-2023), handling inconsistencies, and mitigating hallucinations in numerical data (Hu et al., 2024a; b).
-Current tools such as chain-of-thought (CoT; Wei et al.
-2023), tree-of-thought (ToT; Yao et al.
-2023), Reflexion (Shinn et al., 2023), are designed to provide reasoning traces for LLM decisions; however, their explanations remain ambiguous.
-They lack precise, quantitative insights into key factors and the degree of uncertainty involved.
-Consequently, decision-makers are left with doubts about the reliability of these decisions and how they can be improved.
-There is a pressing need to enhance the verifiability of LLMs in complex decision-making scenarios to ensure their dependability and effectiveness.
-We present DeFine, a new framework designed to build probabilistic factor profiles from transcripts of spoken speech that describe complex scenarios.
-These transcripts are often excessively long, containing ungrammatical sentences, repetitions, hedging, and vagueness (Sawhney et al., 2020; Medya et al., 2022).
-For example, during a quarterly earnings call, a company executive might project a positive revenue outlook to boost investor confidence, despite significant uncertainties surrounding these projections (Mukherjee et al., 2022).
-DeFine constructs a factor profile for each transcript that summarizes essential information into a set of factors and estimates the probabilities of potential outcomes for these factors.
-Moreover, we employ the Bradley-Terry model (Bradley & Terry, 1952) to identify dominant factors and evaluate how these factors collectively impact decision-making.
-Our factor profiles are designed to capture the nuances in spoken transcripts.
-This includes not just what is explicitly stated, but also the implications of details that are omitted, providing a viable method for quantifying uncertainty.
-Our research integrates probabilistic factor profiles with analogical reasoning, a type of reasoning that identifies connections between similar situations to facilitate knowledge transfer from a familiar context to a novel one (Webb et al., 2023; Yasunaga et al., 2024).
-It helps decision-makers draw parallels between current situations and past experiences, thus effectively leveraging historical insights to inform their decisions.
-Instead of relying on traditional text matching, we use factor profiles to retrieve analogous examples, which identifies historical cases with similar levels of uncertainty across key dimensions.
-Analogical reasoning further sets our work apart from traditional Bayesian inference frameworks used in decision-making (Halawi et al., 2024; Lin et al., 2024; Liu et al., 2024).
-The latter often require extensive sampling during inference, which tends to increase inference costs and potentially leads to latency issues.
-The contributions of our research are summarized as follows.
-  • We introduce DeFine, a new framework designed to enhance decision-making in LLMs.
-DeFine utilizes probabilistic factor profiles to quantify uncertainty in complex scenarios, along with analogical reasoning that leverages insights from similar past experiences to guide LLMs in making crucial decisions in novel situations.
-Our framework aims to boost the utility of LLMs by adding a layer of transparency to their decision-making processes.
-  • Our research has produced actionable insights for predicting stock movement trends by analyzing earnings call transcripts.
-These insights enable investors to make informed, data-driven decisions.
-Furthermore, our approach has potential applications beyond finance, including in fields such as medical consultations and political debates (Lehman et al., 2022), where the discussions involve complex issues and the decisions made can have significant consequences.
-[^1] We plan to make the data and code publicly available upon acceptance to facilitate research in this area.
 Figure 1: An excerpt from a typical earnings call transcript and its associated factor profile.
-2 Overview and Factor Analysis
-LLMs have shown great promise in financial services (Reddy et al., 2024; Lee et al., 2024), yet the specifics of their decision-making processes remain largely unexplored.
-In this paper, we investigate how LLMs can guide investment decisions by analyzing earnings call transcripts.
-The transcripts contain a rich mix of textual and numerical data, presenting a unique challenge for LLMs in making rational decisions.
-An earnings call is a teleconference in which the management of a public company discusses its financial results with analysts and investors for a specific reporting period, such as a quarter or a fiscal year.
-[^2] https://en.wikipedia.org/wiki/Earnings_call The transcript generally consists of two sections: the initial prepared remarks from the company’s executives, and a subsequent Q&A session.
-During these calls, executives provide a deep dive into the company’s financials, discuss key performance indicators, and share strategic plans for the future.
-An excerpt from a typical earnings call transcript is shown in Figure 1.
-Our goal is to provide investment recommendations based on earnings call transcripts (ECTs), using a five-tier system: Strong Buy, Buy, Hold, Sell, and Strong Sell.
-We choose this approach over a simple binary classification (Ni et al., 2024) to give a clearer and more nuanced assessment of the investment opportunity.
-We highlight the key drivers in decision-making by identifying a small set of factors from a lengthy transcript (Eigner & Händler, 2024; Feng et al., 2024).
-For example, in an earnings call, discussions of financials such as revenue, expenses, and profit margins can be overwhelming.
-A factor profile helps to distill these discussions into multiple variables, effectively reducing information redundancy and allowing decision-makers to focus on the most impactful factors.
-Crucially, a factor profile offers a comprehensive view on an earnings call.
-If critical elements such as debt levels are not addressed by company executives, they can be marked as ‘unknown or uncertain.’ This contrasts with textual summaries of the transcript (Cho et al., 2021; Khatuya et al., 2024), which may be biased toward the topics emphasized by executives and discussed during the Q&A session.
-2.1 Factor Profile
-[…]
+Large language models are increasingly relied on for decision-making, thanks to their powerful reasoning capabilities OpenAI et al.
+(2024); Anthropic (2025); Kavukcuoglu (2025).
+While research has advanced rapidly in areas such as math, coding, and logical reasoning Bostrom et al.
+(2022); Huang and Chang (2023); Sprague et al.
+(2024); Mondorf and Plank (2024); Li et al.
+(2025); Ren et al.
+(2025), there is growing interest in exploring how LLMs reason through complex, real-world environments to make high-stakes decisions, such as financial investments (Peters, 2024).
+The challenges are compounded when processing long contexts (Krishna et al., 2023; Laban et al., 2024).
+Core issues may include recency bias, hallucinations, numerical inconsistencies, and more (Liu et al., 2023; Hu et al., 2024a, b; Gao et al., 2024).
+While these models are designed to produce reasoning traces, their explanations can remain vague and even unfaithful Chen et al.
+(2025).
+Critically, LLMs still lack precise, quantitative insight into the key factors that lead to their final decisions, as well as mechanisms for incorporating uncertainty into their decision-making.
+In this paper, we present DeFine, a new framework for constructing probabilistic factor profiles from speech transcripts that describe complex scenarios, and leveraging these profiles to enhance decision-making.
+For example, during an earnings call, a company executive might project strong revenue growth to boost investor confidence, despite substantial uncertainty surrounding these projections (Mukherjee et al., 2022).
+As illustrated in Figure 1, DeFine generates a structured factor profile from each transcript, capturing not only what is explicitly stated, but also the implications of what is left unsaid.
+It then uses the BT model (Bradley and Terry, 1952) to identify dominant factors and evaluate how these factors collectively impact decision-making.
+Our research integrates probabilistic factor profiles with analogical reasoning, a type of reasoning that identifies connections between similar situations to facilitate knowledge transfer from a familiar context to a new situation (Webb et al., 2023; Yasunaga et al., 2024).
+Instead of relying on text matching, we use factor profiles to retrieve analogous examples, which are historical cases with similar levels of uncertainty across key dimensions.
+Analogical reasoning further sets our work apart from traditional Bayesian inference models used in decision-making (Halawi et al., 2024; Lin et al., 2024; Liu et al., 2024), which usually lack direct connections to historical cases.
+Our key contributions are as follows.
+  • We introduce DeFine, a modular framework that enhances LLM decision-making in complex scenarios.
+It transforms lengthy speech transcripts into structured factor profiles, which contain key decision factors and their associated uncertainties.
+Analogical reasoning retrieves comparable profiles and draws on similar past cases to make critical decisions.
+DeFine thereby adds a layer of transparency into the decision process.
+  • Our research demonstrates how to effectively extract decision factors from lengthy transcripts and use them to forecast post-earnings stock movements.
+This approach has the potential to generalize to other domains, such as consulting, financial investments, and political debates (Lehman et al., 2022), where discussions are complex and decisions carry significant consequences.
+2 The DeFine Framework
+We investigate how LLMs make critical decisions about post-earnings stock movements using earnings call transcripts (Ni et al., 2024; Peters, 2024).
+An earnings call is a teleconference in which company executives discuss financial results with analysts and investors for a given quarter or fiscal year.
+As shown in Figure 5 in the Appendix, these transcripts typically consist of two parts: prepared remarks from the company’s executives and a subsequent Q&A session.
+During these calls, the executives provide a deep dive into the company’s financials, discuss key performance indicators, and share strategic plans for the future.
+However, discussions of financials, e.g., revenue, expenses, and profit margins, can be overwhelming.
+A factor profile seeks to distills these discussions into multiple variables, allowing decision-makers to focus on the most impactful factors (Eigner and Händler, 2024; Feng et al., 2024).
+Notably, if critical elements such as debt levels are not addressed by company executives, they can be marked as “unknown or uncertain.” This approach contrasts with traditional textual summaries of the transcript (Cho et al., 2021, 2022; Khatuya et al., 2024), which may be biased toward the topics emphasized by executives and discussed during the Q&A session.
+2.1 Constructing Factor Profiles
+Let $X$ be an earnings call transcript used to predict a stock investment decision $Y$ with five outcomes: strong buy, buy, hold, sell, and strong sell.
+We construct a factor profile for each transcript, defining a set of factors =$\mathcal{F}=\{F_{1},F_{2},\dots,F_{n}\}$, where each factor $F_{i}$ has potential outcomes $\{O_{i1},O_{i2},\dots,O_{im}\}$.
+The likelihood of each outcome, given $X$, is modeled by a probabilistic function $P(O_{ij}|X)$.
+These probabilities are inferred using a methodology that optimally integrates textual reasoning with quantitative analysis.
+Thus, each factor outcome’s probability informs the aggregation model that predicts $Y$.
 1.
@@
 Table 1: A curated set of 15 factors for forecasting stock movements following earnings.
-In this study, we focus on a curated set of 15 factors, categorized into three groups: macroeconomic influences (e.g., economic health, market sentiment), company-specific dynamics (e.g., mergers and major acquisitions, product launches), and historical financial metrics (e.g., past earnings, stock prices).
-These factors were carefully selected through an iterative process of querying the LLM for key variables crucial in forecasting stock movements following earnings announcements.
-We intentionally limited our variable set to 15 factors, each with two to three potential outcomes, as detailed below.
-By distilling the analysis to a few significant predictors, our approach balances complexity and performance, while also allowing for future integration of domain-specific factors identified by financial analysts.
-  • Macroeconomic Influences.
-These encompass broad economic factors that affect the entire market or large segments of it.
-This includes the overall economic health, market sentiment, political events, natural disasters and geopolitical issues (Liu et al., 2024).
-Each factor leads to two potential outcomes; for instance, natural disasters might cause a ‘Major Impact’ by disrupting economies and global supply chains, and directly affecting market performance; the ‘Unknown or Uncertain’ outcome reflects the unpredictability of such events.
-  • Company-Specific Dynamics.
-These factors are linked to the internal operations and strategic decisions of individual companies, such as mergers and acquisitions, regulatory changes, financial health, company growth potential, product launches, and issues within the supply chain.
-Each factor can result in one of two potential outcomes.
-For example, a ‘Positive Outlook’ on regulatory changes can open up new business opportunities, whereas ‘Unknown or Uncertain’ could signify regulatory uncertainties that lead to financial challenges.
-  • Historical Financial Metrics.
-Important metrics include historical earnings per share (EPS), revenue trends, and past stock price movements.
-Each factor can result in three outcomes: ‘Bullish’, where metrics like earnings per share, revenue, and stock prices consistently rise, indicating strong financial health; ‘Stable’, characterized by steady movements; ‘Bearish,’ marked by declining financial figures, possibly leading investors to be pessimistic about the company’s future performance.
+Our study focuses on 15 key factors grouped into three categories: macroeconomic influences (e.g., economic health, market sentiment), company-specific dynamics (e.g., mergers and major acquisitions, product launches), and historical financial metrics (e.g., past earnings, stock prices).
+These factors were iteratively selected by querying the LLM for key variables in forecasting stock movements.
+We limit the set to 15 factors, each with two to three possible outcomes, ensuring a balance between complexity and performance while allowing future integration of analyst-identified factors.
+The full list of factors is in the Appendix.
 We make use of the structured output capability of GPT-4o-2024-08-06 to extract factor profiles from earnings call transcripts.
@@
 2.2 Analyzing Key Factors Using the Bradley-Terry Model
-The Bradley-Terry model is a probabilistic framework used for estimating the relative strengths of items based on pairwise comparisons, and the outcome of each comparison indicates which of the two items is ‘better’ in a specific context (Bradley & Terry, 1952).
+The Bradley-Terry model is a probabilistic framework used for estimating the relative strengths of items based on pairwise comparisons, and the outcome of each comparison indicates which of the two items is ‘better’ in a specific context (Bradley and Terry, 1952).
 This model has been widely used for ranking purposes in sports tournaments, LLM preference studies, and other domains where pairwise comparison data is available (Hu et al., 2023; Zhu et al., 2024).
@@
 Table 2: Our dataset includes 11,950 earnings call transcripts from 800+ companies.
+System Recall Prec. F1 Accu. Label Recall Prec. F1
+LLM+CoT+Trans 21.56 33.66 13.52 19.59 Strong Sell 7.32 37.50 12.24
+LLM+CoT+Summ 22.77 16.17 14.12 20.61 Sell 5.56 9.09 6.90
+LLM+CoT+Factors 24.38 28.58 17.26 22.32 Hold 29.84 28.24 29.02
+DeLLMa 38.30 23.14 16.68 22.35 Buy 44.83 18.93 26.62
+DeFine (Ours) 26.15 27.67 23.73 29.64 Strong Buy 43.22 44.56 43.88
+Table 3: (Left) We show the accuracy and macro-averaged F-scores for various systems.
+Our system, DeFine, which combines factor profiles with analogical reasoning, achieves the best performance.
+(Right) DeFine’s performance across five categories: Strong Sell, Sell, Hold, Buy, and Strong Buy.
+Figure 2: A comparison of confusion matrices from the LLM+CoT+Trans, DeLLMa, and DeFine methods.
+While LLM+CoT+Trans and DeLLMa lean towards ‘Buy (B),’ DeFine offers more balanced outcomes across all decision categories, showing notable improvement in ‘Strong Buy (SB),’ ‘Buy (B),’ ‘Hold (H),’ and ‘Sell (S)’ decisions.
 5 Data Collection
@@
 In this section, we evaluate the decision-making performance of various systems, analyze the key factors that influence stock movement predictions, and conduct an analysis of analogical reasoning.
-System Recall Prec. F1 Accu. Label Recall Prec. F1
-LLM+CoT+Trans 21.56 33.66 13.52 19.59 Strong Sell 7.32 37.50 12.24
-LLM+CoT+Summ 22.77 16.17 14.12 20.61 Sell 5.56 9.09 6.90
-LLM+CoT+Factors 24.38 28.58 17.26 22.32 Hold 29.84 28.24 29.02
-DeLLMa (Liu et al., 2024) 38.30 23.14 16.68 22.35 Buy 44.83 18.93 26.62
-DeFine (Ours) 26.15 27.67 23.73 29.64 Strong Buy 43.22 44.56 43.88
-Table 3: (Left) We show the accuracy and macro-averaged F-scores for various systems.
-Our system, DeFine, which combines factor profiles with analogical reasoning, achieves the best performance.
-(Right) DeFine’s performance across five categories: Strong Sell, Sell, Hold, Buy, and Strong Buy.
+Factor (Outcome) Salience - Regulatory changes and legal issues 0.0364 happened (positive outlook) - Natural disasters and other black 0.0360 swan events (major impact) - Political events and government 0.0349 policies (major upheaval) - Geopolitical issues (escalation to 0.0345 conflict) - Supply chain (positive outlook) 0.0322 - Tech innovation (positive outlook) 0.0317 - Historical stock price change (bullish) 0.0316 - Historical EPS (bullish) 0.0315 - Financial health (positive outlook) 0.0311
+Table 4: Influential factors that drive bullish investment decisions in the Consumer Defensive sector, e.g., food and beverage, household products, and grocery stores.
+Factor (Outcome) Salience - Economic health (unknown or uncertain) 0.0362 - Market sentiment and investor 0.0350 psychology (unknown or uncertain) - Company growth (unknown or uncertain) 0.0338 - Supply chain (unknown or uncertain) 0.0326 - Geopolitical issues (escalation to conflict) 0.0322 - Historical revenue (decline) 0.0319 - Historical stock price change (bullish) 0.0318 - Tech innovation (unknown or uncertain) 0.0315 - Natural disasters and other black 0.0315 swan events (major impact) - Political events and government 0.0313 policies (major upheaval)
+Table 5: Factors that drive bullish investment decisions in the Technology sector, including industry leaders such as Apple, Microsoft, Amazon, Google, and Meta.
+Figure 3: We analyze and plot the probability density function (PDF) of positive and negative factor outcomes for different investment decisions.
+Highlighted sections illustrate where the gaps between strong buy (red) and strong sell (blue) decisions are most pronounced.
 6.1 Decision-Making with DeFine
-We test our system, DeFine, against different decision-making strategies: (a) LLM+CoT+Trans: We feed the entire earnings call transcript to the LLM and then use the chain-of-thought to assign a label, with both labels and their interpretations also provided to the LLM.
-(b) LLM+CoT+Summ and LLM+CoT+Factors: These approaches use a summarize-then-predict strategy.
-LLM+CoT+Summ simplifies the transcript into a textual summary, while LLM+CoT+Factors condenses it into a factor profile.
-It is considered a structured summary, unlike the textual summary produced by direct LLM prompting.
-Details on the prompts used for these methods can be found in the Appendix.
+We evaluate our system, DeFine, against various decision-making strategies: (a) LLM+CoT+Trans: The LLM processes the full earnings call transcript, using chain-of-thought reasoning to assign a label with interpretations.
+(b) LLM+CoT+Summ and LLM+CoT+Factors: Both follow a summarize-then-predict approach.
+LLM+CoT+Summ generates a textual summary, while LLM+CoT+Factors condenses information into a structured factor profile.
+Details on the prompts are in the Appendix.
 Our system, DeFine, utilizes analogical reasoning by analyzing five analogous cases identified using KL-divergence as the distance metric.
@@
 In contrast, our factor profile method proves advantageous as it provides a balanced view of both macroeconomic factors and company-specific details, which are essential for rational decision-making.
-Figure 2: A comparison of confusion matrices from the LLM+CoT+Trans, DeLLMa, and DeFine methods.
-While LLM+CoT+Trans and DeLLMa lean towards ‘Buy (B),’ DeFine offers more balanced outcomes across all decision categories, showing notable improvement in ‘Strong Buy (SB),’ ‘Buy (B),’ ‘Hold (H),’ and ‘Sell (S)’ decisions.
 We further analyze DeFine’s performance across five categories: Strong Sell, Sell, Hold, Buy, and Strong Buy.
@@
 According to the F-scores presented in Table 7, all DeFine-BT variants outperform both the random baseline, which assigns investment decisions randomly from five possible labels, and DeLLMa on the test set.
+Tech FS Health CC Ind CS CD Energy RE BM Util Technology (Tech) 15.40 17.99 17.39 12.10 13.15 18.19 25.85 27.09 13.82 22.86 26.67 Financial Services (FS) 15.96 17.96 26.84 7.99 10.21 26.22 15.45 4.80 13.37 21.37 0.00 Healthcare (Health) 16.73 19.80 17.89 21.85 28.46 10.86 20.23 18.73 3.64 43.45 73.33 Consumer Cyclical (CC) 18.14 11.02 19.38 19.39 15.86 9.49 17.70 17.40 12.19 22.22 36.67 Industrials (Ind) 17.02 11.14 14.37 11.24 18.81 15.93 19.48 25.11 3.20 24.44 0.00 Communication Services (CS) 18.61 14.68 18.87 14.03 19.47 33.70 10.71 16.99 11.87 10.26 13.33 Consumer Defensive (CD) 24.91 21.71 19.15 19.89 21.38 2.67 23.09 12.50 9.72 29.52 50.00 Energy 19.49 16.50 23.62 14.25 19.03 8.90 19.03 8.98 12.10 28.79 0.00 Real Estate (RE) 22.86 15.74 16.76 14.08 12.34 4.00 15.61 11.28 12.34 43.18 0.00 Basic Materials (BM) 20.67 13.69 15.26 18.18 29.64 9.52 21.19 17.19 17.10 16.67 37.50 Utilities (Util) 17.82 27.75 23.15 26.25 12.61 25.49 20.63 5.70 18.40 14.29 53.33
+Table 6: The performance of DeFine-BT was evaluated by training it on one financial sector and testing it on another using 100 earnings call transcripts from each of the 11 sectors.
 Among the three variants, DeFine-BT-Cross Sector achieves the highest scores in both F-Score and Accuracy.
@@
 F1 Accu. Random Baseline 18.00 19.11 DeLLMa ((Liu et al., 2024)) 16.68 22.53 DeFine-BT-Same Sector 20.11 22.15 DeFine-BT-Same Company 20.42 23.68 DeFine-BT-Cross Sectors 24.45 27.43
-Table 4: Among the three variants, DeFine-BT-Cross Sector achieves the highest scores, suggesting that considering pairwise comparisons from a diverse range of companies can enhance the predictions of stock movements.
-Factor (Outcome) Salience - Regulatory changes and legal issues 0.0364 happened (positive outlook) - Natural disasters and other black 0.0360 swan events (major impact) - Political events and government 0.0349 policies (major upheaval) - Geopolitical issues (escalation to 0.0345 conflict) - Supply chain (positive outlook) 0.0322 - Tech innovation (positive outlook) 0.0317 - Historical stock price change (bullish) 0.0316 - Historical EPS (bullish) 0.0315 - Financial health (positive outlook) 0.0311
-Table 5: Influential factors and outcomes that drive bullish investment decisions in the Consumer Defensive sector, such as food and beverage, household products, and grocery stores.
-Factor (Outcome) Salience - Economic health (unknown or uncertain) 0.0362 - Market sentiment and investor 0.0350 psychology (unknown or uncertain) - Company growth (unknown or uncertain) 0.0338 - Supply chain (unknown or uncertain) 0.0326 - Geopolitical issues (escalation to conflict) 0.0322 - Historical revenue (decline) 0.0319 - Historical stock price change (bullish) 0.0318 - Tech innovation (unknown or uncertain) 0.0315 - Natural disasters and other black 0.0315 swan events (major impact) - Political events and government 0.0313 policies (major upheaval)
-Table 6: Factors and outcomes that drive bullish investment decisions in the Technology sector, including industry leaders such as Apple, Microsoft, Amazon, Google, and Meta.
-Tables 6 and 6 highlight influential factors impacting investment decisions in the Consumer Defensive and Technology sectors, as identified by the Bradley-Terry model.
+Table 7: DeFine-BT-Cross Sector achieves the highest scores, suggesting that considering pairwise comparisons from a diverse range of companies can enhance the predictions of stock movements.
+Tables 5 and 5 highlight influential factors impacting investment decisions in the Consumer Defensive and Technology sectors, as identified by the Bradley-Terry model.
 In Consumer Defensive, which includes industries like food and beverage, household products, and grocery stores, significant drivers are natural disasters and black swan events, political events and government policies, and geopolitical issues.
@@
 For our system, DeFine, we have opted for $K=5$ to strike a balance between providing enough analogous examples and maintaining a manageable context length for the LLM.
-Figure 3: Analogical reasoning utilizes a number of analogous examples to inform decision-making in LLMs. We adjust $K$ from 3 to 9 and observe its impact on the F-Score.
+Figure 4: Analogical reasoning in LLMs uses analogous examples for decision-making. We vary $K$ from 3 to 9 to assess its impact on the F-Score.
 Moreover, we examine how the most analogous examples influence DeFine’s predictions.
@@
 These results indicate that while DeFine effectively utilizes analogous historical data to inform its predictions, it also critically evaluates the current factor profiles, demonstrating a balanced approach in its decision-making abilities.
-Figure 4: We analyze and plot the probability density function (PDF) of positive and negative factor outcomes for different investment decisions.
-Highlighted sections illustrate where the gaps between strong buy (red) and strong sell (blue) decisions are most pronounced.
 7 Related Work
-Analogical Reasoning.
+Analogical Reasoning
+.
 This type of reasoning identifies connections between similar, though not identical, situations to transfer knowledge from a known context to a new one (Webb et al., 2023; Ozturkler et al., 2023; Yu et al., 2024; Yuan et al., 2024; Sourati et al., 2024; Yasunaga et al., 2024).
@@
 This ability to recognize and use similarities in different situations is important for decision-making.
-While zero-shot analogical reasoning is a desired capability for LLMs, recent studies show they lack the robustness and generality of human analogy-making, as evidenced by counterexamples in tasks such as letter string analogies (Hodel & West, 2024; Lewis & Mitchell, 2024).
+While zero-shot analogical reasoning is a desired capability for LLMs, recent studies show they lack the robustness and generality of human analogy-making, as evidenced by counterexamples in tasks such as letter string analogies (Hodel and West, 2024; Lewis and Mitchell, 2024).
 Musker et al.
@@
 Unlike previous research, our study employs probabilistic factor profiles to model analogical reasoning, grounding our approach in solid mathematical principles.
-Tech FS Health CC Ind CS CD Energy RE BM Util Technology (Tech) 15.40 17.99 17.39 12.10 13.15 18.19 25.85 27.09 13.82 22.86 26.67 Financial Services (FS) 15.96 17.96 26.84 7.99 10.21 26.22 15.45 4.80 13.37 21.37 0.00 Healthcare (Health) 16.73 19.80 17.89 21.85 28.46 10.86 20.23 18.73 3.64 43.45 73.33 Consumer Cyclical (CC) 18.14 11.02 19.38 19.39 15.86 9.49 17.70 17.40 12.19 22.22 36.67 Industrials (Ind) 17.02 11.14 14.37 11.24 18.81 15.93 19.48 25.11 3.20 24.44 0.00 Communication Services (CS) 18.61 14.68 18.87 14.03 19.47 33.70 10.71 16.99 11.87 10.26 13.33 Consumer Defensive (CD) 24.91 21.71 19.15 19.89 21.38 2.67 23.09 12.50 9.72 29.52 50.00 Energy 19.49 16.50 23.62 14.25 19.03 8.90 19.03 8.98 12.10 28.79 0.00 Real Estate (RE) 22.86 15.74 16.76 14.08 12.34 4.00 15.61 11.28 12.34 43.18 0.00 Basic Materials (BM) 20.67 13.69 15.26 18.18 29.64 9.52 21.19 17.19 17.10 16.67 37.50 Utilities (Util) 17.82 27.75 23.15 26.25 12.61 25.49 20.63 5.70 18.40 14.29 53.33
-Table 7: The performance of DeFine-BT was evaluated by training it on one financial sector and testing it on another using 100 earnings call transcripts from each of the 11 sectors.
-LLM Decision-Making under Uncertainty.
+Decision-Making under Uncertainty.
 The use of LLMs in decision-making has surged due to their remarkable ability to reason over complex scenarios (Halawi et al., 2024; Lin et al., 2024; Ye et al., 2024; Band et al., 2024).
@@
 Financial Forecasting.
-Recent advancements in LLMs have revolutionized traditional financial tasks (Keith & Stent, 2019; Sawhney et al., 2020; 2021; Chuang & Yang, 2022; Ang & Lim, 2022; Sang & Bao, 2022; Medya et al., 2022; Wang et al., 2023; Koa et al., 2024; Srivastava et al., 2024).
+Recent advancements in LLMs have revolutionized traditional financial tasks (Keith and Stent, 2019; Sawhney et al., 2020, 2021; Chuang and Yang, 2022; Ang and Lim, 2022; Sang and Bao, 2022; Medya et al., 2022; Wang et al., 2023; Koa et al., 2024; Srivastava et al., 2024).
 Notably, Chen et al.
@@
 The DeFine framework offers a promising avenue for navigating complex data and supporting decision-making processes.
+Acknowledgements
+We are grateful to the reviewers for their insightful feedback, which has helped improve our paper. This research has been partially supported by the NSF CAREER award, #2303655.
 9 Limitations
-The effectiveness of the DeFine framework, as presented in this paper, is predominantly based on controlled experimental conditions.
-While the framework has been designed to enhance decision-making capabilities through the use of probabilistic factor profiles and analogical reasoning, actual outcomes may vary when applied in real-world scenarios.
-Users should be aware that the framework’s performance can be influenced by various external factors including data quality, context-specific nuances, and the dynamic nature of real-world environments.
-We encourage users to consider these variables when implementing and adapting the DeFine approach to ensure its optimal application and to mitigate potential discrepancies between expected and actual results.
-Appendix A Appendix
-[…]
-PromptBox s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark, colback=light!30!white, fonttitle=, fontupper=, title=Constructing a Factor Profile from an Earnings Call Transcript,
-{PromptBox}
-System Message
-You are a financial analyst specializing in earnings call transcripts.
+The DeFine framework, as detailed in our paper, offers a promising approach to enhancing decision-making through its innovative use of probabilistic factor profiles and analogical reasoning.
+Developed under carefully controlled experimental conditions, its potential is noteworthy.
+However, it’s important to acknowledge that its performance might vary in the complexity of real-world settings due to factors such as data quality and specific contextual nuances.
+We encourage users to consider these variables when adapting the DeFine approach to ensure its optimal application and to mitigate potential discrepancies between expected and actual results.
+Appendix A Influential Factors
+  • Macroeconomic Influences.
+These encompass broad economic factors that affect the entire market or large segments of it.
+This includes the overall economic health, market sentiment, political events, natural disasters and geopolitical issues (Liu et al., 2024).
+Each factor leads to two potential outcomes; for instance, natural disasters might cause a ‘Major Impact’ by disrupting economies and global supply chains, and directly affecting market performance; the ‘Unknown or Uncertain’ outcome reflects the unpredictability of such events.
+  • Company-Specific Dynamics.
+These factors are linked to the internal operations and strategic decisions of individual companies, such as mergers and acquisitions, regulatory changes, financial health, company growth potential, product launches, and issues within the supply chain.
+Each factor can result in one of two potential outcomes.
+For example, a ‘Positive Outlook’ on regulatory changes can open up new business opportunities, whereas ‘Unknown or Uncertain’ could signify regulatory uncertainties that lead to financial challenges.
+  • Historical Financial Metrics.
+Important metrics include historical earnings per share (EPS), revenue trends, and past stock price movements.
+Each factor can result in three outcomes: ‘Bullish’, where metrics like earnings per share, revenue, and stock prices consistently rise, indicating strong financial health; ‘Stable’, characterized by steady movements; ‘Bearish,’ marked by declining financial figures, possibly leading investors to be pessimistic about the company’s future performance.
+Transcript Excerpt from Delta Air Lines (DAL) Earnings Call [Prepared Remarks:] >> Operator Good morning, everyone, and welcome to the Delta Air Lines September-quarter 2021 financial results conference call.
+My name is Jen, and I will be your coordinator.
+[Operator instructions] As a reminder, todays call is being recorded.
+I would now like to turn the conference over to Ms.
+Julie Stewart, vice president of investor relations...
+>> Julie Stewart -- Vice President of Investor Relations Thank you, Jen.
+Good morning, everyone, and thanks for joining us for our September-quarter 2021 earnings call.
+Joining us from Atlanta today are CEO, Ed Bastian; our president, Glen Hauenstein; our CFO, Dan Janki.
+And Ed will open the call with an overview of Deltas performance and strategy.
+Glen will provide an update on the revenue environment and our brand momentum, and Dan will discuss cost fleet and our balance sheet.
+Similar to last quarters call, weve scheduled todays call for 90 minutes to make sure we have plenty of time for questions.
+[Operator instructions] After the analyst Q\&A, we will move to our media questions, after which, Ed will provide a brief closing statement.
+Todays discussion contains forward-looking statements that represent our beliefs or expectations about future events.
+All forward-looking statements involve risks and uncertainties that could cause the actual results to differ materially from the forward-looking statements.
+Some of the factors that may cause such differences are described in Deltas SEC filings.
+We also discuss non-GAAP financial measures, and all results exclude special items unless otherwise noted.
+You can find a reconciliation of our non-GAAP measures on the Investor Relations page at ir.delta.com.
+And with that, Ill turn the call over to Ed.
+>> Ed Bastian -- Chief Executive Officer Well, thank you, Julie, and good morning, everyone.
+Appreciate you joining us this morning.
+The September quarter marked another important milestone in our recovery.
+We achieved our first quarterly profit since the start of the pandemic with a pre-tax result of \$216 million and a pre-tax margin of nearly 3% despite still missing one-third of our revenue base compared to the same period in 2019...
+[omitted.] [Questions & Answers:] >> Operator Thank you.
+And well go first to Jamie Baker with J.P.
+Morgan.
+>> Jamie Baker -- J.P.
+Morgan -- Analyst Hey.
+Good morning, everybody.
+First question goes potentially to Glen and Dan.
+So pre-COVID, I had asked Paul about the amount of time that it would typically take Delta to recalibrate the higher fuel prices.
+Im not staring at the transcript, but his estimate at the time was four to six months, which was an improvement from historic levels.
+So my question, I guess, for Glen is whether the booking curve is steep enough right now that you might actually be able to recapture the top line more quickly than that.
+And similarly, for Dan, whether theres anything we should be taking on the cost or operations side that could accelerate the process.
+Im basically just trying to understand whether four to six months is still the right estimate for us to be using.
+>> Glen Hauenstein -- President Well, I would just comment, I think were a bit in uncharted territory here as the recovery continues.
+And while I think it might be difficult in the very short run, despite the fact that the booking curve has moved in a bit, that I would estimate that, that four to six months is about right because we believe that demand and capacity will fall back into a very good equilibrium by next spring which would put you inside that window...
+Figure 5: Transcript Excerpt from Delta Air Lines (DAL) Earnings Call. Adapted from Lu et al. (2025).
+Building a Factor Profile from the Earnings Call Transcript System Message You are a financial analyst specializing in earnings call transcripts.
 You will receive the complete transcript of an earnings call, which includes both the prepared remarks and the Q&A session.
 Your job is to identify the key factors from the transcript and assign probabilities to the potential outcomes of these factors.
-User Message
-Your task is to conduct a comprehensive analysis of the earnings call transcript below.
+User Message Your task is to conduct a comprehensive analysis of the earnings call transcript below.
 Be sure to accurately capture the important factors and estimate the likelihood of each factor resulting in specific outcomes.
@@
 # Prepared Remarks
-Speaker: {Speech…}
+Speaker: {Speech...}
 # Questions and Answers
-Analyst: {Speech…}
+Analyst: {Speech...}
 Please analyze the above earnings call transcript, focusing on the following key factors:
@@
 # Your Output:
-Figure 5: Constructing a Factor Profile from an Earnings Call Transcript
-[…]
-histPrompt s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark,colback=light!30!white, fonttitle=, fontupper=, title=A Prompt to Analyze Trends Based on Historical Financial Metrics,
-[…]
-ARPrompt s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark,colback=light!30!white, fonttitle=, fontupper=, title=Using Analogical Reasoning to Make Investment Decisions,
-{ARPrompt}
-System Message
-You’re a financial analyst who specializes in giving investors buy or sell recommendations by thoroughly analyzing earnings call transcripts.
-User Message
-Here are several example company profiles.
+Figure 6: Building a Factor Profile from the Earnings Call Transcript
+Applying Analogical Reasoning to Investment Decisions System Message You’re a financial analyst who specializes in giving investors buy or sell recommendations by thoroughly analyzing earnings call transcripts.
+User Message Here are several example company profiles.
 Each profile highlights key factors from an earnings call transcript and probabilities for potential outcomes based on those factors.
@@
 ‘idx’ is the index of the most analogous example profile, and ‘recommendation’ should be one of the actions mentioned above for 30 days of trading, and ‘justification’ should clearly explain your recommendation using the strategy you learned from the selected example company profile.
-Figure 6: Using Analogical Reasoning to Make Investment Decisions
-[…]
-CoTPrompt s O !htbp halign=left, floatplacement=#2, IfBooleanTF=#1float,width=float, colframe=dark,colback=light!30!white, fonttitle=, fontupper=, title=A Prompt that Uses Chain-of-Thought to Make Investment Decisions,
-{CoTPrompt}
-System Message
-You’re a financial analyst spcializing in giving investors buy or sell recommendations by thoroughly analyzing earnings call transcripts.
-User Message
-Based on your analysis of the earnings call for {Company Name} held on {Announcement Date}, decide on the most likely analyst recommendation for the next 30 days from these options:
- - Action 1: strong buy: The stock price will increase by more than 5%
+Figure 7: Applying Analogical Reasoning to Investment Decisions.
+Chain-of-Thought Prompt for Investment Decisions System Message You’re a financial analyst spcializing in giving investors buy or sell recommendations by thoroughly analyzing earnings call transcripts.
+User Message Based on your analysis of the earnings call for {Company Name} held on {Announcement Date}, decide on the most likely analyst recommendation for the next 30 days from these options:
+- Action 1: strong buy: The stock price will increase by more than 5%
  - Action 2: buy: The stock price will increase by 2% to 5%
@@
 ‘Thoughts’ should be your detailed reasoning steps, ‘recommendation’ should be one of the actions mentioned above for 30 days trading, ‘Justification’ should clearly explain your recommendation using the strategy you learned from the selected example company profile.
-Figure 7: A Prompt that Uses Chain-of-Thought to Make Investment Decisions
-{histPrompt}
-System Message
-You are a financial analyst specializing in historical data analysis, including stock prices, earnings per share (EPS), and revenue.
+Figure 8: Chain-of-Thought Prompt for Soliciting Investment Decisions.
+Prompt for Analyzing Trends in Historical Financial Data System Message You are a financial analyst specializing in historical data analysis, including stock prices, earnings per share (EPS), and revenue.
 Your goal is to assess the likelihood of different market trends based on past data.
-User Message
-The potential outcomes to consider are: {bullish, stable, and bearish}.
+User Message The potential outcomes to consider are: {bullish, stable, and bearish}.
 For each outcome, please assign a likelihood level from the following options: {very likely, likely, somewhat likely, somewhat unlikely, unlikely, very unlikely}.
@@
  2023-08-01  195.46
-{… until the date of the earnings announcement.}
+{...until the date of the earnings announcement.}
 Please analyze this historical data and provide the likelihood of each outcome in JSON format.
@@
 # Your Output:
-Figure 8: A Prompt to Analyze Trends Based on Historical Financial Metrics
+Figure 9: Prompt for Analyzing Trends in Historical Financial Data.